# SATO: Stable Text-to-Motion Framework

### v2

@@ Abstract.
Is the Text to Motion model robust? Recent advancements in Text to Motion models primarily stem from more accurate predictions of specific actions. However, the text modality typically relies solely on pre-trained Contrastive Language-Image Pretraining (CLIP) models. Our research has uncovered a significant issue with the text-to-motion model: its predictions often exhibit inconsistent outputs, resulting in vastly different or even incorrect poses when presented with semantically similar or identical text inputs. In this paper, we undertake an analysis to elucidate the underlying causes of this instability, establishing a clear link between the unpredictability of model outputs and the erratic attention patterns of the text encoder module. Consequently, we introduce a formal framework aimed at addressing this issue, which we term the Stable Text-to-Motion Framework (SATO). SATO consists of three modules, each dedicated to stable attention, stable prediction, and maintaining a balance between accuracy and robustness trade-off. We present a methodology for constructing an SATO that satisfies the stability of attention and prediction. To verify the stability of the model, we introduced a new textual synonym perturbation dataset based on HumanML3D and KIT-ML. Results show that SATO is significantly more stable against synonyms and other slight perturbations while keeping its high accuracy performance.
@@
*Equal contribution.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 1. Introduction
The Text-to-Motion (T2M) model signifies a groundbreaking and swiftly advancing paradigm with immense potential across various domains, such as video games, the metaverse, and virtual/augmented reality environments. This innovative approach, as evidenced by research contributions from (Zhang et al., 2023b; Ghosh et al., 2021; Guo et al., 2022b, 2023; Tevet et al., 2022b; Zhang et al., 2023a) revolves around generating motion data directly from textual descriptions, thereby simplifying the overall process and mitigating associated time and cost overheads.
@@
Figure 4. Visual results on user testing. SATO (T2M-GPT) refers to fine-tuning based on T2M-GPT to create SATO. Below each action sequence is the corresponding motion caption. The red color text represents the top-k attention weight words. It can be seen that the perturbation of the caption can lead to changes in the attention of the text, which can lead to catastrophic errors in the generative model. SATO has demonstrated superior stability to other models both in terms of attention and motion prediction. More visual results are provided in Supplementary Material Section C and on the project website.
@@ 2.1. Text-conditioned human motion generation
Text-conditioned human action generation aims to generate 3D human actions based on textual descriptions. Recent mainstream work can be divided into two categories, namely VQ-VAE-based methods and diffusion models. VQ-VAE (van den Oord et al., 2018; Esser et al., 2021; Williams et al., 2020; Ramesh et al., 2021; Ao et al., 2022; Dhariwal et al., 2020) has achieved excellent performance in multi-modal generation tasks. ACTOR (Petrovich et al., 2021) proposes a Transformer-based VAE for generating motion from predefined action categories. TEMOS (Petrovich et al., 2022) introduces an additional text encoder based on ACTOR for generating different action sequences based on text descriptions, but mainly focusing on short sentences. Guo et al. (Guo et al., 2022a) propose an autoregressive conditional VAE conditioned on the generated frame and text features, and proposed to predict actions based on the length of the text. TEACH (Athanasiou et al., 2022) is based on TEMOS, which generates temporal motion combinations from a series of natural language descriptions and extends space for long action combinations. TM2T (Guo et al., 2022b) considers not only text-to-motion tasks, but also motion-to-text tasks, and the joint training of these two tasks will be improved. T2M-GPT (Zhang et al., 2023b) quantizes motion clips into discrete markers and then uses a converter to generate subsequent markers. The emerging diffusion models are also changing the field of motion generation. MDM (Tevet et al., 2022b) uses a Transformer Encoder as the main body of prediction samples. MotionDiffuse (Zhang et al., 2022) uses the DDPM architecture to generate realistic and diverse motion.
However, whether it is the diffusion-based method the VQ-VAE-based method, or even previous work such as MotionCLIP (Tevet et al., 2022a), the structure is based on the CLIP encoder. Although the work of TEMOS, TEACH, and Guo et al. considered the sequence length of the text and the time and space issues, they did not take into account the diversification of text raised by users. When the text is subject to slight perturbation, the model may exhibit inconsistent outputs, even leading to catastrophic errors in motion, which is a common and severe problem with these past methods. Therefore, based on these issues, this paper is the first work to consider the diversity of user-proposed texts and the first work based on the stable framework in the field of text-conditional human action generation. In this paper, we propose SATO so that the text generation results can still show strong robustness when encountering synonyms or other slight replacements or interference.
2.2. Stable Text-to-Motion
@@
For the stabilization of input vector perturbations, some work has been done on stabilizing the output pattern of the model from various perspectives. Reconstructing the perturbed text with the actual input text can improve the the robustness of the text model (Sato et al., 2018), but does not guarantee the model’s attention similarity before and after the perturbation. Cansu et al. (Sen et al., 2020) analyze human and machine attention to the text. However, they fail to analyze the consistency of the descriptions before and after perturbation. Compared to the text after the perturbation, the model is more inclined to use ”unfamiliar vocabulary” in human comprehension. As the ”unfamiliar vocabulary” increases in a description, it interferes with the comprehension of the text. Shunsuke et al. (Kitada and Iyatomi, 2021) enhance the text embedding stability using adversarial learning but do not analyze the consistency between old and new attention, making it difficult to ensure that textual attention can have consistent results for different descriptive scenarios under the same semantics. Yin et al. (Yin et al., 2022) employ an adversarial robustness approach to enhance the stability of NLP models. However, these techniques are designed for ex-post interpretation of model predictions and thus cannot be applied to enhance attentional stability in the prediction phase. Different from text or visual stable attention, for the multimedia text-to-motion domain, we not only need to consider the distribution of attention weights during text embedding but also focus on the coordination performance between text and motion generation. Therefore, we need to pay attention to the consistency of textual attention before and after the description perturbation. And we also pay attention to the local importance and overall compatibility of semantic weights, to avoid the repeated generation of the emphasized part of the description, thus ignoring the coherence of the whole action.
@@ 3.1. Preliminaries
VQ-VAE Based Text-to-Motion Model. Our objective is to produce a 3D human pose sequence $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}]$, where $\mathbf{x}_{t}\in\mathbb{R}^{d}$, guided by a textual description $\mathbf{C}=[c_{1},c_{2},...,c_{l}]$, where $T$ represents the number of frames and $d$ denotes the dimension of the motion feature. Here, $\mathbf{c}_{i}$ represents the $i^{th}$ word in the sentence, and $l$ is the length of the sentence. The process begins with extracting a text embedding $\mathbf{c_{e}}$ from input text using CLIP. Subsequently, a transformer model predicts the distribution of possible next indices $p(S_{i}|\mathbf{c_{e}},S_{<i})$ based on the text embedding $\mathbf{c_{e}}$ and previous indices $S_{<i}$, where $S_{i}$ represents the index of the next element at position $i$ in the sequence. These predicted indices are then mapped to corresponding entries in the learned codebook, yielding latent code representations $\mathbf{\hat{z}_{i}}$ (van den Oord et al., 2018). Finally, the decoder network decodes these codebook entries into motion sequences $\mathbf{X_{pred}}$. The optimization objective aims to maximize the log-likelihood of the data distribution. This is achieved by denoting the likelihood of the full sequence as $p(S|\mathbf{c_{e}})=\prod_{i=1}^{|S|}p(S_{i}|\mathbf{c_{e}},S_{<i})$ and directly maximizing it: $\mathcal{L}_{\text{trans}}=\mathbb{E}_{S\sim p(S)}[-\log p(S|\mathbf{c_{e}})]$ (Zhang et al., 2023b; Guo et al., 2023), facilitating the generation of motion sequences from input text.
@@ 3.2. Problem Formulation
(1) $V_{k}(\mathbf{x},\mathbf{x}^{\prime})=\frac{1}{k\cdot|T_{k}(\mathbf{x})\cap T_{k}(\mathbf{x}^{\prime})|}.$
@@
Balancing Accuracy and Robustness. Accuracy and robustness are naturally in a trade-off relationship (Zhang et al., 2019; Raghunathan et al., 2020). Our objective is to bolster stability while minimizing the decline in model accuracy, thereby mitigating catastrophic errors arising from input perturbations. Consequently, we require a mechanism to uphold the model’s performance concerning the original input.
@@
    (Prediction Robustness) $D_{1}(\tilde{y}(\mathbf{x},\tilde{\mathbf{\omega}})$,$\tilde{y}(x,\tilde{\mathbf{\omega}}+\boldsymbol{\boldsymbol{\rho_{1}}}))\leq\gamma_{1}$, for some $\|\boldsymbol{\boldsymbol{\rho}}_{1}\|\leq R_{1},\gamma_{1}\geq 0$
@@
    (Top-$k$ Attention Robustness) $V_{k}(\tilde{\mathbf{\omega}},\tilde{\mathbf{\omega}}+\boldsymbol{\boldsymbol{\rho}_{2}})\geq\beta$ ,for some $1\geq\beta\geq 0$, $\|\boldsymbol{\boldsymbol{\rho}}_{2}\|\leq R_{2}$;
@@ 3.3. Stable Text-to-Motion Framework
MotionDiffuse (Zhang et al., 2022) arXiv2022 0.630±.001 2.435±.067 1.549±.032 0.491±.001 0.681±.001 0.782±.001 3.113±.001 9.410±.049
MDM (Tevet et al., 2022b) ICLR2023 0.544±.044 3.251±.071 2.015±.027 -- -- 0.611±.007 5.566±.027 9.559±.086
@@
MotionDiffuse arXiv2022 1.954±.062 5.737±.172 2.496±.106 0.417±.004 0.621±.004 0.739±.004 2.958±.005 11.100±.143
MDM ICLR2023 0.497±.021 3.564±.894 2.331±.032 -- -- 0.396±.004 9.191±.022 10.847±.109
@@
MoMask CVPR2024 0.204±.011 2.570±.092 2.234±.101 0.433±.007 0.656±.005 0.781±.005 2.779±.022 2.779±.022
SATO (T2M-GPT) -- 0.513±.006 0.581±.005 0.137±.002 0.410±.011 0.619±.005 0.736±.005 3.123±.034 10.889±.066
@@
Table 3. Human evaluation and cross-dataset results on the original or perturbed text. ”Excellent” means completely meets the semantic, with smooth and correct expression; ”Good” means generally generates well with minor details; ”Fair” means contains errors in details but is overall correct; ”Poor” means overall incorrect; ”Very poor” means motions and text cannot be matched at all. We believe that Excellent, Good, and Fair represent correctly generated postures, while the other two represent errors. Preference indicates human preference for the compared motions. The cross-dataset evaluation result is that the model is trained on the HumanML3D dataset, with text from KIT-ML used for testing.
@@ 4. Experiments
Notely, to measure the stability of the model, we will use three different $FID$ input calculation methods: (1) $FID$: the distribution distance between the motion generated from the original text and real motion. (2) $FID_{P}$: the distribution distance between the motion generated from text after paraphrasing and real motion. (3) $FID_{D}$: the distribution distance between the motion generated from the original text and the motion generated from the text after paraphrasing. Moreover, we also employ human evaluation for cross-dataset evaluation to further analyze the performance and stability of the model. For assessing the stability of model attention, we propose using JSD. A smaller JSD value indicates greater stability of attention under perturbation. More details about the evaluation metrics are provided in Supplementary Material Section B.
@@ 4.2. Comparison with SOTA
Stability. It is worth noting that all other models perform poorly on the perturbed dataset, with $FID_{P}$ significantly greater than $FID$, indicating that diverse representations of perturbations are fatal to the performance of these models. On $FID_{P}$, SATO(T2M-GPT) significantly reduced by 1.599 and 2.175 on HumanML3D and KIT-ML respectively, while SATO(Momask) decreased by 0.899 and 2.311 respectively. Similarly, there was a significant increase in $FID_{D}$, with SATO(T2M-GPT) decreasing by 1.422 and 2.757 respectively, and SATO(MoMask) decreasing by 1.058 and 2.178 respectively. This suggests that SATO yields similar predictive results on both perturbed and original datasets, indicating stronger stability. We can also observe a significant reduction in the fluctuation of our model on the $FID,FID_{P},FID_{D}$ metrics, which also reflects the stability of our approach in predictions. We further investigated the impact of our approach on the attention JSD metric. Our method exhibits stability in attention, as evidenced by experiments and visualizations provided in the supplementary material.
Accuracy. Although our model experiences a slight decrease in $FID$ and R-precision, we would like to point out that previous work has shown that a small decrease in $FID$ does not necessarily imply a decrease in generation quality (Jayasumana et al., 2024). Our visualization results corroborate this point, and we have additional examples from supplementary material and anonymous website to further substantiate this perspective. Furthermore, compared to the significant improvement in stability measured by the $FID$ metric, the slight decrease in $FID$ on the original text can be almost disregarded. Our visualizations and additional human evaluation experiments also demonstrate that the quality of text generated by our model on both original and perturbed text is superior to the original model. This means that our model can generate higher-quality motion sequence outputs in practical applications, with a lower likelihood of catastrophic errors occurring when presented with a broader range of textual inputs.
@@
Figure 5. Model stability evaluation under different perturbations. The x-axis represents texts with varying degrees of perturbation, while the left y-axis denotes $FID_{D}$ and the right y-axis represents $FID_{P}$. It can be observed that across all levels of perturbation, SATO (T2M-GPT) consistently outperforms T2M-GPT in terms of stability metrics. Even when subjected to significant perturbation, our model maintains excellent stability.

### v3

@@ Abstract.
Is the Text to Motion model robust? Recent advancements in Text to Motion models primarily stem from more accurate predictions of specific actions. However, the text modality typically relies solely on pre-trained Contrastive Language-Image Pretraining (CLIP) models. Our research has uncovered a significant issue with the text-to-motion model: its predictions often exhibit inconsistent outputs, resulting in vastly different or even incorrect poses when presented with semantically similar or identical text inputs. In this paper, we undertake an analysis to elucidate the underlying causes of this instability, establishing a clear link between the unpredictability of model outputs and the erratic attention patterns of the text encoder module. Consequently, we introduce a formal framework aimed at addressing this issue, which we term the Stable Text-to-Motion Framework (SATO). SATO consists of three modules, each dedicated to stable attention, stable prediction, and maintaining a balance between accuracy and robustness trade-off. We present a methodology for constructing an SATO that satisfies the stability of attention and prediction. To verify the stability of the model, we introduced a new textual synonym perturbation dataset based on HumanML3D and KIT-ML. Results show that SATO is significantly more stable against synonyms and other slight perturbations while keeping its high accuracy performance. Codes and models are released at https://github.com/sato-team/Stable-Text-to-Motion-Framework
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681034isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Computer vision
@@ 1. Introduction
The Text-to-Motion (T2M) model signifies a groundbreaking and swiftly advancing paradigm with immense potential across various domains, such as video games, the metaverse, and virtual or augmented reality environments. This innovative approach, as evidenced by research contributions from (Zhang et al., 2023b; Ghosh et al., 2021; Guo et al., 2022b, 2023; Tevet et al., 2022b; Zhang et al., 2023a) revolves around generating motion data directly from textual descriptions, thereby simplifying the overall process and mitigating associated time and cost overheads.
@@
Figure 4. Visual results on user testing. SATO (T2M-GPT) refers to fine-tuning based on T2M-GPT to create SATO. Below each action sequence is the corresponding motion caption. The bold text represents the top-k attention weight words. It can be seen that the perturbation of the caption can lead to changes in the attention of the text, which can lead to catastrophic errors in the generative model. SATO has demonstrated superior stability to other models both in terms of attention and motion prediction.
@@ 2.1. Text-conditioned human motion generation
Text-conditioned human motion generation aims to generate 3D human actions based on textual descriptions. Recent mainstream work can be divided into two categories, namely VAE-based methods and diffusion models. VAE (van den Oord et al., 2018; Esser et al., 2021; Williams et al., 2020; Ramesh et al., 2021; Ao et al., 2022; Dhariwal et al., 2020) has achieved excellent performance in multi-modal generation tasks. ACTOR (Petrovich et al., 2021) proposes a Transformer-based VAE for generating motion from predefined action categories. TEMOS (Petrovich et al., 2022) introduces an additional text encoder based on ACTOR for generating different motion sequences based on text descriptions, but mainly focusing on short sentences. Guo et al. (Guo et al., 2022a) propose an autoregressive conditional VAE conditioned on the generated frame and text features, and proposed to predict motions based on the length of the text. TEACH (Athanasiou et al., 2022) is based on TEMOS, which generates temporal motion combinations from a series of natural language descriptions and extends space for long motion combinations. TM2T (Guo et al., 2022b) considers not only text-to-motion tasks, but also motion-to-text tasks, and the joint training of these two tasks will be improved. T2M-GPT (Zhang et al., 2023b) quantizes motion clips into discrete markers and then uses a converter to generate subsequent markers. However, whether it is the diffusion-based method (Tevet et al., 2022b; Zhang et al., 2023a; Liu et al., 2024a) the VAE-based method, or even previous work such as MotionCLIP (Tevet et al., 2022a), the structure is based on the CLIP encoder (Lin et al., 2024). When the text is subject to slight perturbation, the model may exhibit inconsistent outputs, even leading to catastrophic errors in motion, which is a common and severe problem with these past methods.
2.2. Robustness of text conditioned motion generation
@@
For the stabilization of input vector perturbations, some work has been done on stabilizing the output pattern of the model from various perspectives. Reconstructing the perturbed text with the actual input text can improve the the robustness of the text model (Sato et al., 2018), but does not guarantee the model’s attention similarity before and after the perturbation. Shunsuke et al. (Kitada and Iyatomi, 2021) enhance the text embedding stability using adversarial learning but do not analyze the consistency between old and new attention, making it difficult to ensure that textual attention can have consistent results for different descriptive scenarios under the same semantics. In text-to-motion field, Wang et al. (Wang et al., 2023) introduced a multi-layer graph neural network to learn the local features and correlations of tokens in action descriptions. Jin et al. (Jin et al., 2023) constructed a hierarchical semantic graph to enrich the details of action generation while maintaining the overall coherence of the actions.
However, for multidimensional descriptions that the text embedding model has not encountered, the inherent challenge lies in the static text model’s difficulty in reading characteristic description information, making it hard for the model to align the semantic features of descriptions across different dimensions (Liu et al., 2024b). Therefore, we need to pay attention to the consistency of textual attention before and after the description perturbation. And we also pay attention to the local importance and overall compatibility of semantic weights, to avoid the repeated generation of the emphasized part of the description, thus ignoring the coherence of the whole action.
@@ 3.1. Preliminaries
VQ-VAE Based Text-to-Motion Model. Our objective is to produce a 3D human pose sequence $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}]$, where $\mathbf{x}_{t}\in\mathbb{R}^{d}$, guided by a textual description $\mathbf{C}=[c_{1},c_{2},...,c_{l}]$, where $T$ represents the number of frames and $d$ denotes the dimension of the motion feature. Here, $\mathbf{c}_{i}$ represents the $i^{th}$ word in the sentence, and $l$ is the length of the sentence. The process begins with extracting a text embedding $\mathbf{c_{e}}$ from input text using CLIP. Subsequently, a transformer model predicts the distribution of possible next indices $p(S_{i}|\mathbf{c_{e}},S_{<i})$ based on the text embedding $\mathbf{c_{e}}$ and previous indices $S_{<i}$, where $S_{i}$ represents the index of the next element at position $i$ in the sequence. These predicted indices are then mapped to corresponding entries in the learned codebook, yielding latent code representations $\mathbf{\hat{z}_{i}}$ (van den Oord et al., 2018). Finally, the decoder network decodes these codebook entries into motion sequences $\mathbf{X_{pred}}$. The optimization objective aims to maximize the log-likelihood of the data distribution. This is achieved by denoting the likelihood of the full sequence as $p(S|\mathbf{c_{e}})=\prod_{i=1}^{|S|}p(S_{i}|\mathbf{c_{e}},S_{<i})$ and directly maximizing it: $\mathcal{L}_{\text{trans}}=\mathbb{E}_{S\sim p(S)}[-\log p(S|\mathbf{c_{e}})]$ (Zhang et al., 2023b; Guo et al., 2023; Cai et al., 2024), facilitating the generation of motion sequences from input text.
@@ 3.2. Problem Formulation
(1) $V_{k}(\mathbf{x},\mathbf{x}^{\prime})=\frac{1}{k\cdot|T_{k}(\mathbf{x})\cap T_{k}(\mathbf{x}^{\prime})|}$
@@
Balancing Accuracy and Robustness. Accuracy and robustness are naturally in a trade-off relationship (Zhang et al., 2019; Raghunathan et al., 2020; Lai et al., 2024; Hu et al., 2024; Li et al., 2024). Our objective is to bolster stability while minimizing the decline in model accuracy, thereby mitigating catastrophic errors arising from input perturbations. Consequently, we require a mechanism to uphold the model’s performance concerning the original input.
@@
    (Prediction Robustness) $D_{1}(\tilde{y}(\mathbf{x},\tilde{\mathbf{\omega}})$,$\tilde{y}(x,\tilde{\mathbf{\omega}}+\boldsymbol{\boldsymbol{\rho_{1}}}))\leq\gamma_{1}$, for some $\|\boldsymbol{\boldsymbol{\rho}}_{1}\|\leq R_{1},\gamma_{1}\geq 0$.
@@
    (Top-$k$ Attention Robustness) $V_{k}(\tilde{\mathbf{\omega}},\tilde{\mathbf{\omega}}+\boldsymbol{\boldsymbol{\rho}_{2}})\geq\beta$, for some $1\geq\beta\geq 0$, $\|\boldsymbol{\boldsymbol{\rho}}_{2}\|\leq R_{2}$.
@@ 3.3. Stable Text-to-Motion Framework
MotionDiffuse (Zhang et al., 2024) TPAMI2024 0.630±.001 2.435±.067 1.549±.032 0.491±.001 0.681±.001 0.782±.001 3.113±.001 9.410±.049
MDM (Tevet et al., 2022b) ICLR2023 0.544±.044 6.283±.154 5.428±.160 -- -- 0.611±.007 5.566±.027 9.559±.086
@@
MotionDiffuse TPAMI2024 1.954±.062 5.737±.172 2.496±.106 0.417±.004 0.621±.004 0.739±.004 2.958±.005 11.100±.143
MDM ICLR2023 0.497±.021 7.631±.217 8.334±.226 -- -- 0.396±.004 9.191±.022 10.847±.109
@@
MoMask CVPR2024 0.204±.011 2.570±.092 2.234±.101 0.433±.007 0.656±.005 0.781±.005 2.779±.022 10.734±.008
SATO (T2M-GPT) -- 0.706±.041 0.679±.041 0.090±.041 0.380±.005 0.587±.006 0.709±.005 3.203±.024 10.860±.107
@@
Table 3. Human evaluation and cross-dataset results on the original or perturbed text. The cross-dataset evaluation result is that the model is trained on the HumanML3D dataset, with text from KIT-ML used for testing.
@@ 4. Experiments
Notely, to measure the stability of the model, we will use three different $FID$ input calculation methods: (1) $FID$: the distribution distance between the motion generated from the original text and real motion. (2) $FID_{P}$: the distribution distance between the motion generated from text after paraphrasing and real motion. (3) $FID_{D}$: the distribution distance between the motion generated from the original text and the motion generated from the text after paraphrasing. Moreover, we also employ human evaluation for cross-dataset evaluation to further analyze the performance and stability of the model. For assessing the stability of model attention, we propose using JSD. A smaller JSD value indicates greater stability of attention under perturbation.
@@ 4.2. Comparison with SOTA
Stability. It is worth noting that all other models perform poorly on the perturbed dataset, with $FID_{P}$ significantly greater than $FID$, indicating that diverse representations of perturbations are fatal to the performance of these models. On $FID_{P}$, SATO (T2M-GPT) significantly reduced by 1.599 and 2.077 on HumanML3D and KIT-ML respectively, while SATO (Momask) decreased by 0.899 and 2.311 respectively. Similarly, there was a significant increase in $FID_{D}$, with SATO (T2M-GPT) decreasing by 1.422 and 2.804 respectively, and SATO (MoMask) decreasing by 1.058 and 2.178 respectively. This suggests that SATO yields similar predictive results on both perturbed and original datasets, indicating stronger stability. We can also observe a significant reduction in the fluctuation of our model on the $FID,FID_{P},FID_{D}$ metrics, which also reflects the stability of our approach in predictions. We further investigated the impact of our approach on the attention JSD metric. Our method exhibits stability in attention, as evidenced by experiments and visualizations provided in the supplementary material.
Accuracy. Although our model experiences a slight decrease in $FID$ and R-precision, we would like to point out that previous work has shown that a small decrease in $FID$ does not necessarily imply a decrease in generation quality (Jayasumana et al., 2024). Furthermore, compared to the significant improvement in stability measured by the $FID$ metric, the slight decrease in $FID$ on the original text can be almost disregarded. Our visualizations and additional human evaluation experiments also demonstrate that the quality of text generated by our model on both original and perturbed text is superior to the original model. This means that
textbfour model can generate higher-quality motion sequence outputs in practical applications, with a lower likelihood of catastrophic errors occurring when presented with a broader range of textual inputs.
@@
Figure 5. Model stability evaluation under different perturbations. It can be observed that across all levels of perturbation, SATO (T2M-GPT) consistently outperforms T2M-GPT in terms of stability metrics. Even when subjected to significant perturbation, our model maintains excellent stability.